OpenStudy (anonymous):

@jigglypuff314 @ganeshie8

OpenStudy (anonymous):

Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation.

jigglypuff314 (jigglypuff314):

distribute in 4(7x - 1) what do you get?

OpenStudy (anonymous):

um, idk.. cant think of it

OpenStudy (anonymous):

24x

jigglypuff314 (jigglypuff314):

use the distributive property a(b-c) = ab - ac so 4(7x-1) = ?

OpenStudy (anonymous):

4*7 - 4*1 28 - 4 24

jigglypuff314 (jigglypuff314):

close! :) remember that you can't add or subtract terms that have different amounts of x so 28x - 4 can't be subtracted it would just be left like that but now do you see the difference in what "Max" did wrong? :)

OpenStudy (anonymous):

never mind, i kept reading the equation wrong, haha... ok so the mistake that he subtracted 28x - 4,

jigglypuff314 (jigglypuff314):

mistake was http://prntscr.com/3u2jqh

OpenStudy (anonymous):

OH oh my god i am so stupid, i really read that thing wrong, haha

OpenStudy (anonymous):

thanks, i get it now

jigglypuff314 (jigglypuff314):

I'm glad I could help :)